zarkam21 (zarkam21):

Solve the system of equations. Enter your answer as an ordered pair. 3x+5y=48 -3x+5y=12

OpenStudy (anonymous):

just add them and what do you get ?

OpenStudy (anonymous):

notice that when adding, the 3x's cancel out.....now you can solve for y

zarkam21 (zarkam21):

12?

OpenStudy (anonymous):

3x + 5y = 48 -3x + 5y = 12 ------------add 10y = 60 -- now divide both sides by 10 y = 6 now sub 6 in for y back in either of the original equations and find x questions ?

zarkam21 (zarkam21):

6?

OpenStudy (anonymous):

we know that y = 6 so we will sub that back in for y in one of the original equations 3x + 5y = 48 3x + 5(6) = 48 3x + 30 = 48 3x = 48 - 30 3x = 18 x = 6 now we can check. -3x + 5y = 12 -3(6) + 5(6) = 12 -18 + 30 = 12 12 = 12 (correct) so the solution is (6,6) questions ?

zarkam21 (zarkam21):

Nope

OpenStudy (anonymous):

good to hear :)